# Mon CHÈRI $\heartsuit$ Adapting Capability Hardware Enhanced RISC with Conditional Capabilities

### v1

@@ III-B High-Level System Architecture
The conditional capability-enhanced CHERI processor (Figure 3(a)) is modified to support the conditional capability instructions (➃, § IV-A), and its instruction pipeline is augmented with logic to encode and decode capabilities containing an additional operation bound (➄, § III-A). Modified ➅ store and ➆ load logic updates and checks the operation bounds when executing store and load instructions, respectively. Finally, ➇ the pipeline performs writebacks of updated capabilities in register operands after the operations that update the operation bounds. Code capability checks for the Program Counter Capability (PCC) are done in a dedicated bounds check block ➈ (needed for Write-before-Execute, Write-before-Execute-Only, and Execute-Once), while all data CPs use a general-purpose bounds check block ➉.
IV Implementation
@@
In this section, we present Mon CHÈRI, our implementation of CPs for the CHERI-RISC-V architecture. We implemented two versions of Mon CHÈRI: 1. a Mon CHÈRI software model for the QEMU-system-CHERI128 full-system emulator, and 2. a Mon CHÈRI softcore based on the CHERI-Flute64 processor intellectual property (IP). The ISA extension that forms the basis for Mon CHÈRI is described in § IV-A.
@@
In addition, we extended the existing CHERI-LLVM toolchain [19] with support for the Mon CHÈRI ISA extension and Write-before-Read CP instrumentation for stack-allocated variables. Our changes to CHERI-LLVM are described in § IV-C. Finally, we added run-time Write-before-Read CP support for heap-allocated memory in the CheriFreeRTOS [1] and Two-Level Segregated Fit (TLSF) [15] memory allocators. Our changes to the allocators are described in § IV-D.
IV-A ISA extension for CHERI-RISC-V
TABLE II: CSetOpBOunds family of instructions.
Mnemonic Conditional permission
csetwbrbound Write-before-Read
csetwbxbound Write-before-Execute
csetrobound Write-before-Read-Only
csetxobound Write-before-Execute-Only
csetwtbound Write-Once
csetrtbound Read-Once
csetxtbound Execute-Once
IV-B Adding operation bounds to compressed capabilities
@@
Figure 4: Layouts of 128-bit CHERI Capabilities without (4(a)) and with (4(b)) compressed operation bound. The labels indicate the fields for permissions ($p$), flag ($f$), CP control bits ($p_{op}$), object type (otype), internal exponent ($I_{E}$), compressed top ($T$) including high part of exponent ($T_{E}$), compressed base ($B$) including low part of exponent ($B_{E}$), and cursor consisting of the pointer address ($a$) in 4(a), the address and compressed operation bound ($O$, $O_{E}$) in 4(b).
CSetOpBounds instructions.
The RISC-V ISA is extensible through portions of the instruction encoding space reserved for ISA extensions. We add a family of CSetOpBounds instructions to CHERI-RISC-V to allow setting an initial, or updating an already set, value for a capability’s operation bound. Our CSetOpBounds instructions require two operands: an existing capability and a length operand. When invoked, CSetOpBounds sets the operation bound to the range $\left[b,o\right]$ where $b$ is the “base” encoded as part of the conventional capability bound, and $o$ is the “operation top” encoded into the operation bound (see § IV-B). CSetOpBounds does not allow programs to increase the operation bounds while they can still can invoke CSetOpBounds to decrease the operation bounds of a capability. The operational bounds set by CSetOpBounds, in any mode, are restricted to be within the original capability bounds.
Additionally, CSetOpBounds must identify the CPs which operation bound to set. Lacking free operands in the instruction encodings available to us, we chose to encode the CP information into the CSetOpBounds opcode. We place the CSetOpBounds family of instructions into the existing CHERIv9 [79] non-standard Xcheri extension in the custom-2/rv128 opcode space. The full set of CSetOpBounds instructions is shown in Table II (§ IV-A).
Conditional permissions.
We assign new conditional permissions control bits in the previously unused portion of the CHERI capability representation (see Figure 4(b)). The CHERI-Flute64 IP reserves 12-bits for hardware permissions, 4-bits for user permissions, thus leaving 3-bits unused. The CP control bit for the Write-before-Read CP, PERM_WRITE_BEFORE_READ_DISABLE, is by default set to 1 in newly created capabilities. In this default state, the enforcement of Write-before-Read is disabled for the capability that can freely be used to dereference into uninitialized memory within its bounds. Invoking csetwbropbound with the capability as an operand sets the PERM_WRITE_BEFORE_READ_DISABLE control bit to 0 and activates the enforcement of the Write-before-Read CP.
Processor pipeline changes.
For bounds checks, it is important not to increase the critical path within the processor. The execute stage of the pipeline (see Figure 3(b)) has two stages; the CHERI implementation sets up data capability bounds check in the first stage (where single cycle ALU operations are performed), and performs the actual check in the second stage (long-latency operations such as memory access). As operations bounds are very similar, they use a similar procedures where checks are done in parallel to the checks in the general-purpose bounds checker to not incur further latency. For a store, the first stage checks if the requested address is higher than the current operation bound. If the address to be written is higher but adjacent to the existing operation bound, the second stage increases the operation bound by the written amount. In this case, the updated operation bound is committed and written back to register in the writeback stage. For a load, the first stage checks whether the address falls within the operation bounds. In case of any bound violation, a CHERI protection violation is raised during the second stage by triggering a hardware exception.
@@
The CHERI ISA introduced compression for its capability representation in ISAv6 [81]. The current compression scheme, CHERI Concentrate[84] was introduced in ISAv7 [80]. The reason for employing compression for the in-memory representation is simple: the raw 256-bit capability representation quadruples the size of the native 64-bit pointers, increasing cache footprint and memory bandwidth utilization. Increasing the architectural size of capabilities to make space for additional operation bounds is, therefore, not practical. Fortunately, most modern 64-bit operating systems use only a subset of the 64-bits in virtual addresses. For example, 64-bit RISC-V Linux uses 48-bit virtual addresses [28], while 32-bit operating systems, such FreeRTOS, use only 32-bits while running on 64-bit hardware. We utilize these unused bits in the baseline architecture address store to conditional capability operation bounds, allowing up to two 16-bit operation bounds, or one 32-bit operation bound, to be stored in a 128-bit capability.
@@
Masking a portion of address bits in cases where operating systems don’t fully utilize 64-bit addresses has precedents both in conventional processor architectures as well as in the CHERI design. Examples of such pointer masking include Arm’s Top Byte Ignore (TBI) [71], Intel Linear Address Masking (LAM) [33], and AMD Upper Address Ignore (UAI) [3] features, as well as the alternative compression formats in Appendix E of the CHERI ISAv9 [79]. To encode operation bounds in the address, we implemented pointer masking in Mon CHÈRI by introducing an address mask when a register holds a conditional capability. CHERI-RISC-V uses a merged registers file, which means general-purpose registers can hold either a 64-bit integer or a 128-bit capability. Unlike previous proposals for RISC-V pointer masking [48], the Mon CHÈRI address mask is only applied to conditional capabilities. Conventional CHERI capabilities and regular pointers are unaffected.
@@
For the 128-bit ISA variant, the compressed representation relies on redundancy between the three 64-bit virtual addresses the raw 256-bit capability is comprised of: the base ($b$), top ($t$), and the address ($a$), and exchanges stronger alignment requirements (proportional to object size) for a more compact representation.
Figure 4(a) shows the capability format for CHERI Concentrate. $B$ and $T$ encode the $b$ and $t$ bounds in one of two formats depending on the internal exponent ($I_{E}$) bit: if $I_{E}=1$ then an exponent ($E$) is stored in the lower three bits of $B$ and $T$ ($B_{E}$ and $T_{E}$) reducing their precision by three bits. $E$ determines the position at which $B$ and $T$ are inserted into $a$ to obtain $b$ and $t$. Otherwise ($I_{E}=0$, $E=0$) the full width of $b$ and $t$ are used. Their width is determined by an encoding parameter: mantissa width ($MW$) that determines the precision of the decoded bounds. The CHERI ISAv9 uses $MW=14$ for 128-bit capabilities. However, $t$ is further compressed by two bits based on the observation that the top two bits of $t$ can be derived from the equation $T=B+L$ where the most significant bit of $L$ ($L_{msb}$) is known from the values of $I_{E}$ and $E$ and a carry bit is implied if $T\left[11:0\right]<B\left[11:0\right]$ since $t$ is known to be larger than $b$.
@@
Mon CHÈRI compressed capability representation.
Figure 4(b) shows the capability format for Mon CHÈRI. We encode the operation bound, operation top ($o$) in two fields in the most significant 16-bits of the cursor: $O\left[13:3\right]$ (11 bits) and $O_{E}$ (5 bits) which are freed by limiting $a_{top}=a\left[47:E+MW\right]$. When $I_{E}=0$, $o$ is stored identically to $b$ with its lowest three bits $O\left[2:0\right]$ derived from the most significant bits of $O_{E}$ ($O_{E}\left[4:2\right]$). When $I_{E}=1$, up to five bits from $O_{E}\left[E+2:0\right]$ are used to store the least significant bits of $o$. Our current implementation limits $O_{E}$ to five bits limiting $o$ to an $I_{E}$ of at most $2$. We discuss methods to alleviate this limitation in § VII. Expanding $o$ to full 48 bits happens similarly to $b$ with its own correction $c_{o}$. § VI in § -A illustrates the changes relative to CHERI Concentrate compression.
IV-C Compiler Support for CHERI-LLVM
@@
LABEL:lst:ir shows an excerpt of the LLVM IR of one of the functions in LABEL:lst:attributes after CP instrumentation (see ➀ in Figure 3). The CHERI-LLVM compiler introduces capability bounds on stack variables through an LLVM IR-level pass, CheriBoundAllocas, executed just before instruction selection. This pass inserts a llvm.cheri.cap.bounds.set intrinsic for every alloca (stack allocation) instruction (❶ in LABEL:lst:ir). Consequently, all uses of alloca instructions are replaced with llvm.cheri.cap.bounds.set which are then replaced with CHERI csetbounds instructions. We extended the existing LLVM pass to insert new llvm.cheri.cap.op.bounds.set intrinsics (❷ in LABEL:lst:ir) under these conditions:
  • A
    To all stack variables if the application is compiled with the -cheri-write-before-read option.
  • B
    To all variables in functions with the __attribute__(("writebeforeread")) annotation.
  • C
    To variables with the __writebeforeread annotation.
The llvm.cheri.cap.op.bounds.set intrinsic is replaced with csetwbrbound instructions during instruction selection.
The original CheriBoundAllocas performs static program analysis to check if the alloca instructions fall within the conventional CHERI bounds and avoids introducing csetbounds for stack allocations that are within bounds. We disable this optimization for variables that receive llvm.cheri.cap.bounds.set as it cannot statically guarantee that stores to Write-before-Read variables fall within the operation bounds. In § VII, we discuss the possibility of leveraging compiler-based static analysis to also optimize Write-before-Read variables.
@@
In static single-assignment (SSA) form used by the LLVM IR, a variable must be assigned exactly once, and each use of the variable must refer to that single assignment. When dealing with multiple assignments to the same variable, the compiler must transform the assignments into a form that maintains the SSA property. This involves introducing new variables for each assignment to ensure that each variable has a single definition. When the same variable is accessed through different control-flow paths, the compiler introduces a special kind of assignment called a phi ($\Phi$) function. The phi function merges different values coming from different control-flow paths.
During the register allocation phase, a compiler maps the variables in SSA form to the registers of the processor. Usually, the register pressure (i.e., the number of simultaneously live variables at an instruction) in a program is much higher than the number of available registers. Thus, the compiler has to generate so-called spill code that moves the contents of the variables between memory and registers.
The analysis that decides how to split the live range of a variable at the right places for efficient spilling is called live-range splitting. Live-range splitting can make use of the fact that the same variable may, at times, be available in multiple registers simultaneously. This causes challenges for Write-before-Read CP instrumentation as conditional capabilities carry with them state in the form of the operation bound, which is implicitly updated as a result of store operations on the capability. Optimization algorithms do not take this property into account and can cause conditions in which two instances of the same capability exist simultaneously but contain a different representation of its operation bound. This can occur for two reasons: 1. register spilling, and 2. register forking” when conditional capability state being divided across two instances becomes forked across separate registers. Register spilling becomes problematic when a copy of a capability is stored in memory, becomes stale over time, and is later restored.
@@
To avoid forked capabilities and spilled capabilities from going stale, we introduce a new store linearization pass (➁ in Figure 3) that runs after optimizations on the SSA representation, and that manipulates the SSA representation of the program to ensure register allocation can maintain a single, canonical instance of a conditional capability though which all stores occur. Performing store linearization after optimizations allows for robustness across different optimization levels.
Maintaining conditional capability liveness.
@@
The LLVM back end handles instruction selection and register allocation. To ensure the live-range splitting algorithm in the back end avoids spilling a canonical conditional capability, causing it to grow stale, the store linearization inserts placeholder function calls for every store operand in the IR. This placeholder function is effectively a no-op, taking a conditional capability operand used for the store as an input and returning the same operand. However, this introduces an explicit data dependency between the input and output operand, indicating to the live-range spilling algorithm that the conditional capability has changed.
@@
Array indexing via the LLVM IR GetElementPtr (GEP) instruction requires special handling: GEP takes a pointer operand and offset and returns another pointer (in SSA-form) pointing at the specific offset. In this case, the store linearization inserts the placeholder function call for the GEP operand, not the pointer used for indexing. We found that, in practice, the back end will generate code that uses integer-relative store (s[bhwd]) instructions, taking a capability operand and an offset using the GEP input instead of the output operand. Therefore, store linearization targets the original operand to extend its liveness. Finally, the store linearization pass recursively replaces the operands used for store operations with the operands returned by the placeholder function.
@@
When a value in SSA-form escapes from a code block, it must be updated when accessed along alternative control-flow paths. The reg2mem transformation [42] part of LLVM is designed to handle escaped values by: 1) allocating stack memory for each escaped SSA register, 2) storing SSA values to these allocated memory locations before exiting a block, and 3) reloading the value values from memory upon entering a new block. This ensures that all SSA values remain consistent and up-to-date across different code paths. Store linearization uses a slightly modified version of reg2mem to update conditional capabilities across code paths, similarly to how IR-level phi functions merge such values in SSA-form. The original reg2mem pass tracks output operands from GEP instruction as escaped values. The special handling of array indexes explained above allows us to ignore output operand GEP instruction and avoid populating them in memory.
Similarly, conditional capabilities explicitly stored in memory by the program must be updated after they have been used for store operations, even in straight-line code. Store linearization performs a recursive check to determine if the operand pointer used in a store is itself stored in memory. If that is the case, an additional store operation is inserted after the operation bound has been updated to update the capability in memory. LABEL:lst:64bsink and LABEL:lst:64bsinkir in Appendix illustrate the transformation that occurs in this case.
IV-D Write-before-Read Support for Memory Allocators

### v2

@@ 3.3 High-Level System Architecture
The conditional capability-enhanced CHERI processor (Figure 3(a)) is modified to support the conditional capability instructions (➃, § 4.1), and its instruction pipeline is augmented with logic to encode and decode capabilities containing an additional operation bound (➄, § 3.2). Modified ➅ store and ➆ load logic updates and checks the operation bounds when executing store and load instructions, respectively. Finally, ➇ the pipeline performs writebacks of updated capabilities in register operands after the operations that update the operation bounds. Code capability checks for the Program Counter Capability (PCC) are done in a dedicated bounds check block ➈ (needed for Write-before-Execute, Write-before-Execute-Only , and Execute-Once), while all data CPs use a general-purpose bounds check block ➉.
4 Mon CHÉRI Implementation
TABLE II: Architectural changes to CHERI-RISC-V by conditional permission. The CSetOpBounds variants column shows the corresponding CSetOpBounds instruction mnemonic for each CP. The Pipeline changes column shows the impact on the Load, Store, and Execute logic inside the processor. A ○ indicates the operation is unaffected, a ◐ indicates the operation is augmented with an OB check, a ◑ indicates the operation is augmented with a writeback that, under certain conditions, updates the conditional capability OB. Finally, a ● indicates the operation is augmented with both.
Conditional Permission CSetOpBounds variants Pipeline changes
Load Store Execute
Write-before-Read csetwbrbound
Write-before-Execute csetwbxbound
Write-before-Read-Only csetrobound
Write-before-Execute-Only csetxobound
Write-Once csetwtbound
Read-Once csetrtbound
Execute-Once csetxtbound
@@ 4 Mon CHÉRI Implementation
In this section, we present Mon CHÉRI, our implementation of CPs for the CHERI-RISC-V architecture. We implemented two versions of Mon CHÉRI: 1. a Mon CHÉRI software model for the QEMU-system-CHERI128 full-system emulator, and 2. a Mon CHÉRI softcore based on the CHERI-Flute64 processor intellectual property (IP). The Mon CHÉRI ISA extension is described in § 4.1.
@@
In addition, we extended the existing CHERI-LLVM toolchain [30] with support for the Mon CHÉRI ISA extension and Write-before-Read CP instrumentation for stack-allocated variables (§ 4.3). Finally, we added run-time Write-before-Read CP support for heap-allocated memory in the CheriFreeRTOS [31] and Two-Level Segregated Fit (TLSF) [32] memory allocators (§ 4.4).
4.1 Mon CHÉRI Extension for CHERI-RISC-V
CSetOpBounds instructions.
The RISC-V ISA is extensible through portions of the instruction encoding space reserved for ISA extensions. We add a family of CSetOpBounds instructions to CHERI-RISC-V to the existing CHERIv9 [16] non-standard Xcheri extension in the custom-2/rv128 opcode space to allow setting an initial, or updating an already set, value for a capability’s operation bound. Our CSetOpBounds instructions require two operands: an existing capability and a length operand. When invoked, CSetOpBounds sets the operation bound to the range $\left[b,o\right]$ where $b$ is the “base” encoded as part of the conventional capability bound, and $o$ is the “operation top” encoded into the operation bound (see § 4.2). CSetOpBounds does not allow programs to increase the operation bounds, while they are allowed to invoke CSetOpBounds to decrease the operation bounds of a capability. The operational bounds set by CSetOpBounds are restricted to be within the original capability bounds.
Additionally, CSetOpBounds must identify the CPs which operation bound to set. Lacking free operands in the instruction encodings available in Xcheri, we chose to encode the CP information into the CSetOpBounds opcode. The CSetOpBounds instruction variants are shown in Table II.
Conditional permissions.
We assign new conditional permissions control bits from the CHERI capability representation ($p_{op}$ in Figure 4(b)). The CHERI-Flute64 IP reserves 12-bits for hardware-defined permissions, i.e., those authorizing load, store, etc., and 4-bits for software-defined permissions, which interpretation are left open by the CHERIv9 specification. For the purposes of prototyping CPs, we utilize the four available software-defined bits, treating them as a 4-bit integer that enumerates one of 16 possible, mutually exclusive permission states. Five of those states correspond to the CPs in table II, ten modes are left unused, and one mode, the zero value, corresponds to the default state in which CP enforcement is disabled. In this default state, protections such as Write-before-Read are inactive, allowing uninitialized memory access within the capability’s bounds. Invoking csetwbropbound on a capability enables Write-before-Read enforcement for that capability.
Processor pipeline changes.
To avoid increasing processor latency, bounds checks are carefully structured within the execute instruction block of the pipeline (see Figure 3(b)), which has two stages. In the first stage, where single-cycle arithmetic logic unit (ALU) operations are performed, the CHERI implementation checks the conditional permissions and initiates the data capability bounds check, while the actual bounds check occurs in the second stage (used for longer-latency operations like memory access). OB checks run in parallel with the general-purpose bounds checking to minimize their latency. For stores and loads, the first stage checks if the conditional permission allows the operation, and if the target address results in an update of the current operation bound; if it does and is adjacent, the second stage extends the bound to cover it, with the update committed in the writeback stage. Table II shows how the OB checks and writeback are used by different CPs. Any violation raises a CHERI protection exception in the second stage.
@@ 4.1 Mon CHÉRI Extension for CHERI-RISC-V
Figure 4: Layouts of 128-bit CHERI Capabilities without (4(a)) and with (4(b)) compressed operation bound. The labels indicate the fields for software-defined and hardware-defined permissions ($p_{sw}$ and $p_{hw}$ respectively), CP control bits ($p_{op}$), flag ($f$), object type (otype), internal exponent ($I_{E}$), compressed top ($T$) including high part of exponent ($T_{E}$), compressed base ($B$) including low part of exponent ($B_{E}$), and cursor consisting of the pointer address ($a$) in 4(a), the address and compressed operation bound ($O$, $O_{E}$) in 4(b).
Avoiding data hazards.
Conditional Capabilities create dependencies between instructions that are typically independent. This is because stores using conditional capabilities update the OB in the capability operand, while stores on conventional capabilities do not. As a result, when a store using a conditional capability is immediately followed by a load or another store using the same capability, a data hazard occurs: the latter operation requires the updated OB at the ALU stage for the OB check but is not available in the operand register until the writeback stage completes.
To solve this, we implemented a bypass—a new data path within the pipeline—that forwards the updated OB from the memory access or writeback stage to the ALU stage. This bypass activates when the processor detects that the next instruction in the pipeline operates on the same conditional capability. This avoids the need to stall the processor or reorder instructions during compilation and is compatible with conventional RISC-V instruction ordering. Appendix § A.1 shows an example of a data hazard that is avoided by the bypass.
4.2 Adding Operation Bounds to Capabilities
@@
The CHERI ISA introduced capability compression in ISAv6 [33], with the current CHERI Concentrate[19] compression scheme introduced in ISAv7 [34]. Compression reduces the in-memory size of capabilities, which would otherwise occupy 256 bits, quadrupling the space needed for native 64-bit pointers, and increasing cache footprint and memory bandwidth requirements. Expanding capabilities further to include additional operation bounds is impractical. However, most modern 64-bit operating systems (OSs) use only part of the 64-bit virtual address space. For instance, RISC-V Linux uses 48-bit addresses [35], and 32-bit OSs like FreeRTOS use 32 bits even on 64-bit hardware. We leverage these unused bits to store an additional 16-bit OB.
@@
Masking a portion of the address in cases where OSs do not fully utilize 64-bit addresses has precedents in both conventional processors, such as in Arm’s Top Byte Ignore (TBI) [36], Intel Linear Address Masking (LAM) [37], and AMD Upper Address Ignore (UAI) [38] features, as well in the CHERI design as alternative compression formats described in Appendix E of the CHERI ISAv9 [16]. To encode operation bounds, Mon CHÉRI applies an address mask when a register holds a conditional capability. Since CHERI-RISC-V uses a merged register file, general-purpose registers can hold either a 64-bit integer or a 128-bit capability. Unlike prior RISC-V pointer masking proposals [39], Mon CHÉRI’s address mask is limited to conditional capabilities, leaving conventional CHERI capabilities and regular pointers unchanged.
@@
A raw 256-bit capability is comprised of three virtual addresses: base ($b$), top ($t$), and address ($a$). The 128-bit representation of capabilities utilizes the redundancy between the three addresses and stronger alignment requirements (proportional to object size) for a more compact representation.
Figure 4(a) shows the capability format for CHERI Concentrate. $B$ and $T$ encode the $b$ and $t$ bounds in one of two formats depending on the internal exponent ($I_{E}$) bit: if $I_{E}=1$ then an exponent ($E$) is stored in the lower three bits of $B$ and $T$ ($B_{E}$ and $T_{E}$) reducing their precision by three bits. $E$ determines the position at which $B$ and $T$ are inserted into $a$ to obtain $b$ and $t$. Otherwise ($I_{E}=0$, $E=0$) the full width of $b$ and $t$ are used. Their width is determined by an encoding parameter: mantissa width ($MW$) that determines the precision of the decoded bounds. The CHERI ISAv9 uses $MW=14$ for 128-bit capabilities. However, $t$ is further compressed by two bits as the top two bits of $t$ can be derived from the equation $T=B+L$ where the most significant bit of $L$ ($L_{msb}$) is known from the values of $I_{E}$ and $E$ and a carry bit is implied if $T\left[11:0\right]<B\left[11:0\right]$ since $t$ is known to be larger than $b$.
@@
Mon CHÉRI compressed capability representation.
Figure 4(b) shows the capability format for Mon CHÉRI. We encode the operation bound, operation top ($o$) in two fields in the most significant 16-bits of the cursor: $O\left[13:3\right]$ (11 bits) and $O_{E}$ (5 bits) which are freed by limiting $a_{top}=a\left[47:E+MW\right]$. When $I_{E}=0$, $o$ is stored identically to $b$ with its lowest three bits $O\left[2:0\right]$ derived from the most significant bits of $O_{E}$ ($O_{E}\left[4:2\right]$). When $I_{E}=1$, up to five bits from $O_{E}\left[E+2:0\right]$ are used to store the least significant bits of $o$. Our current implementation limits $O_{E}$ to five bits, limiting $o$ to an $I_{E}$ of at most $2$. We discuss methods to alleviate this limitation in § 7. Expanding $o$ to full 48 bits happens similarly to $b$ with its own correction $c_{o}$. Figure 6 in § A.2 illustrates the changes relative to CHERI Concentrate compression.
4.3 Mon CHÉRI Support for CHERI-LLVM
@@
LABEL:lst:ir shows an excerpt of the LLVM IR of one of the functions in LABEL:lst:attributes after CP instrumentation (see ➀ in Figure 3). The CHERI-LLVM compiler adds capability bounds to stack variables in an IR-level compiler pass, CheriBoundAllocas. This pass replaces every IR stack allocation instruction (alloca, ❶ in LABEL:lst:ir) with a llvm .cheri .cap .bounds .set intrinsic. These are then replaced with CHERI csetbounds instructions by the CHERI-RISC-V backend. We extended this pass to add llvm .cheri .cap .op .bounds .set intrinsics  (❷ inLABEL:lst:ir)for variables that are either annotated with __writebeforeread ( C in fig. 3(a)), in functions annotated with __attribute__(("writebeforeread")) ( B), or all stack variables, when compiling with the -cheri-write-before-read option ( A). The llvm .cheri .cap .op .bounds .set intrinsic is replaced with csetwbrbound instructions by the backend.
Optimizing Write-before-Read.
The CheriBoundAllocas pass checks if alloca instructions fall within the original capability bounds and omits llvm .cheri .cap .bounds .set in those cases. We disable this optimization for variables that receive llvm .cheri .cap .op .bounds .set since it does not guarantee stores to Write-before-Read variables fall within the OBs. However, not all variables require run-time CP checks. For Write-before-Read, we optimize arrays and scalar variables allocated in function entry blocks by using a simple, non-heuristic analysis that checks if they are fully initialized before being accessed. The analysis inspects each store in the function’s first basic block and verifies whether they are preceded by loads to the same allocation. For arrays, we track initialization with a vector, checking store instructions for the base pointer or specific indices, ensuring each load is preceded by a store through dominance analysis[40]; if a load occurs before a store, the index is marked uninitialized in the vector. For scalars, we verify whether a store targets the variable, and each load is dominated by a corresponding store. Variables shown to be initialized do not receive the Write-before-Read CP and, therefore, can be checked by CheriBoundAllocas for capability-bound optimization. In § 7, we discuss the possibility of using heuristic static analysis to optimize Write-before-Read variables further.
@@
LLVM IR uses static single-assignment (SSA) form, where each variable has exactly one assignment. The compiler creates new IR variables to maintain this rule when a variable has multiple assignments. If a variable is accessed through different control-flow paths, a phi ($\Phi$) function is introduced to merge the values from these paths.
During register allocation, the compiler maps variables to processor registers. However, due to register pressure—the number of live variables exceeding the number of available registers—–the compiler generates spill code that moves variable contents between memory and registers. Live-range splitting optimizes when variables are spilled and can make use of the fact that the same variable may, at times, be available in multiple registers simultaneously. However, when a conditional capability is stored in multiple registers, the state of its operation bound may become inconsistent across instances. This can occur due to: 1. register spilling, or 2. register forking, where the conditional capability state differs between registers. Inconsistencies arise when a duplicated capability is stored in memory, becomes stale as its copy’s operation bound is updated, and is later restored.
@@
To address this, we introduce a store linearization pass, which runs after SSA optimizations and ensures a single, canonical conditional capability instance is maintained across stores. Performing store linearization after optimizations guarantees robustness across optimization levels.
@@
Store linearization inserts placeholder function calls for every store operand in the IR to prevent live-range splitting from spilling a canonical conditional capability, causing it to grow stale. This placeholder is an identity function—effectively a no-op—taking a conditional capability operand and returning it unchanged. This, however, creates a data dependency between the input and output operand, signaling to the live-range splitting algorithm that the conditional capability has changed since its use in the store.
@@
Array indexing via the LLVM IR GetElementPtr (GEP) instruction requires special handling. GEP takes a pointer and an offset, returning a new pointer (in SSA-form) pointing at the specified offset in the array. Store linearization inserts a placeholder function for the GEP operand, not the indexing pointer. In practice, the CHERI-RISC-V backend generates code using integer-relative store instructions (s[bhwd]), where the GEP input is used as a capability operand with an offset. Therefore, the linearization targets the original operand to extend its liveness.
Finally, the pass recursively replaces any variables holding conditional capabilities used store operands with those returned by the corresponding placeholder functions.
@@
When a value in SSA form escapes a code block, it must be updated when accessed along different control-flow paths. The reg2mem transformation [41] in LLVM handles escaped values by: 1) allocating stack memory for each escaped SSA register, 2) storing SSA values in this memory before exiting a block, and 3) reloading values upon entering a new block. This ensures consistency across code paths. Store linearization modifies reg2mem to update conditional capabilities across control-flow paths, similar to how phi functions merge SSA values. One optimization available to store linearization stems from the original reg2mem also tracking output operands from the GEP instruction as escaped values. The special handling of array indexing allows us to avoid storing output operands from GEP in memory.
Similarly, conditional capabilities stored explicitly in memory must be updated after being used for store operations, even in straight-line code. Store linearization recursively checks if the operand used in a store is itself stored in memory. If so, an additional store operation is inserted to update the capability stored in memory. Listings LABEL:lst:64bsink and LABEL:lst:64bsinkir in § A.3 illustrate this transformation.
4.4 Mon CHÉRI Support for Memory Allocators